OpenStudy (anonymous):

If one bulb in a string of holiday lights fails to work, the whole string will not light. If each bulb in a set has a 99.5% chance of working, what is the maximum number of lights that can be strung together with at least a 90% chance of the whole string lighting? the key is 21

OpenStudy (anonymous):

Key is 21 wth?

OpenStudy (anonymous):

just 21

OpenStudy (dumbcow):

each bulb has probability of 0.995 of working each bulb is independent The combined probability of a string of bulbs working is the product of all the individual bulb probabilities. 0.995*0.995 is the probability that a string of 2 bulbs work and so on \[0.995^{n} = 0.9\] \[n = \frac{\log 0.9}{\log 0.995} = 21.02\]

OpenStudy (anonymous):

I can do this in my head ROFL you do not need all of that

OpenStudy (anonymous):

good job, dumbcow

OpenStudy (dumbcow):

no problem...you can divide logs in you head???

OpenStudy (dumbcow):

i dunno how else to do it

OpenStudy (anonymous):

I have to use caculator

OpenStudy (anonymous):

thank , can't do in my head , I'm slow

OpenStudy (anonymous):

OK so in my head i think of a black bord then i do the steps in my head small step by small step then when i am about to move on i take note of the # i am at, it helps a lot I would not recommend this as you can do some stupid things xD but i think it is faster

OpenStudy (anonymous):

Gopeder was probably just pulling your leg. It's a little hard to believe that someone who has problems with 2 system equations can work out logarithmic equations in their head. Also, the answer has to be round *up* to 22, not down to 21. As you already know, you can't have .02 lightbulbs.